1. A test charge of +2µC is located 4m to the east of a -5µC charge.(k= 9xNm2/C2) A) Find the electric force felt by the test charge. B) Find the electric field at that location.

2. Calculate the electric field felt by a positive test charge located half way between a charge of +2C and a charge of -2C, that are 3m apart


1
Expert's answer
2021-04-13T09:36:18-0400

"(1) \space Q_1=+2\mu C\\Q_2=-5\mu C"

"(A) \space F=\dfrac{kQ_1Q_2}{r^2}=\dfrac{9\times10^9\times2\times5\times10^{-12}}{16}=5.625\times10^{-3}\space N"


"(B)\space E=\dfrac{kQ_2}{r^2}=\dfrac{9\times 10^9\times 5\times 10^{-6}}{16}=2.8125\times10^{3}\space N/C"


"(2)\space Q_1=+2C"

"Q_2=-2 C"


"E_1=\dfrac{9\times10^9\times2}{(1.5)^2}=8\times10^9\space N/C"

"E_2=\dfrac{9\times10^9\times2}{(1.5)^2}=8\times10^9\space N/C"

"E_{res}=E_1+E_2=16\times10^9\space N/C(towards\space left)"